OpenStudy (yumeybitaa):

Let f(x) = x2 - 16. Find f-1(x). (x2 means x to the 2 )

OpenStudy (yumeybitaa):

(f-1--> means f to the -1 )

OpenStudy (anonymous):

\[f(x)=x^2-16\] and you want \[f^{-1}(x)\] right?

OpenStudy (yumeybitaa):

esactly

OpenStudy (anonymous):

ok good tell the idiot who wrote this question that since \[y=x^2-16\] is a parabola, it is evidently not a one to one function so for example \[f(5)=f(-5)=9\] that there is no function \[f^{-1}(x)\]

OpenStudy (anonymous):

on the other hand, we can still solve \[x=y^2-16\] for \(y\) we get \[x+16=y^2\\ \pm\sqrt{x+16}=y\]

OpenStudy (anonymous):

Are you given a domain and range?

OpenStudy (anonymous):

so i have no idea what you are supposed to put perhaps \[f^{-1}(x)=\pm\sqrt{x+16}\]

OpenStudy (anonymous):

the \(\pm\) part tells you it is not a function

OpenStudy (yumeybitaa):

yea i think that's the answer

OpenStudy (anonymous):

ask whoever wrote this question if they know what a function is, and if so, how you can have a function that is either plus or minus depending on your mood the answer is you cannot

OpenStudy (yumeybitaa):

well this question is from my exam of flvs

OpenStudy (anonymous):

let me make a guess FLVS??

OpenStudy (anonymous):

how did i know?

OpenStudy (anonymous):

i keep asking for question id numbers from people who ask these questions' but no one ever wants to tell me

OpenStudy (anonymous):

there questions are the most poorly written i have ever seen written by people who have no idea about math it is a sin you should insist on going to a real school, unless of course the schools in florida are actually even worse

OpenStudy (anonymous):

*their

OpenStudy (yumeybitaa):

i left school because of bullying .....but hey can you help me with another one ?

OpenStudy (anonymous):

sure

OpenStudy (yumeybitaa):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x+24 -40x + 31 -40x+64 -40x+70

OpenStudy (anonymous):

ok we work from the inside out \[f(g(x))=f(10x-6)\] since \[\large f(\heartsuit)=-4\heartsuit+7\] you get \[f(10x-6)=-4(10x+6)-7\] now some algebra

OpenStudy (anonymous):

ok i messed that up hold on

OpenStudy (yumeybitaa):

okay :)

OpenStudy (anonymous):

\[f(10x-6)=-4(10x-6)+7\]thats better

OpenStudy (anonymous):

now algebra \[-40x+24+7\\ -40x+31\]

OpenStudy (anonymous):

go with B

OpenStudy (anonymous):

got more or is that it?

OpenStudy (yumeybitaa):

if you wanna keep helping me i have 2 more

OpenStudy (anonymous):

sure no problem

OpenStudy (yumeybitaa):

Let f(x) = 12 over the quantity of 4 x + 2. Find f(-1).

OpenStudy (anonymous):

\[y=\frac{12}{4x+2}\]

OpenStudy (anonymous):

switch x and y \[x=\frac{12}{4y+2}\] and solve for \(y\)

OpenStudy (anonymous):

\[x=\frac{12}{4y+2}\\ x(4y+2)=12\] \[4xy+2x=12\\ 4xy=12-2x\\ y=\frac{12-2x}{4x}\]

OpenStudy (anonymous):

other in answers include \[f^{-1}(x)=\frac{6-x}{2x}\]

OpenStudy (yumeybitaa):

this is a little bit complicated

OpenStudy (anonymous):

did you see an answer that looked like \[f^{-1}(x)=\frac{6-x}{2x}\]

OpenStudy (yumeybitaa):

yes

OpenStudy (yumeybitaa):

-6 -2 2 6 I'm guessing is 6

OpenStudy (anonymous):

oh wait, all you need is \(f(-1)\)?

OpenStudy (anonymous):

\[f(-1)=\frac{12}{4\times (-1)+2}=\frac{12}{-2}=-6\]

OpenStudy (anonymous):

go with \(-6\)

OpenStudy (yumeybitaa):

phh okay thank you

OpenStudy (anonymous):

next?

OpenStudy (yumeybitaa):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (anonymous):

this one is quick solve \[2x+2=4\] for \(x\) in two steps

OpenStudy (anonymous):

\[2x+2=4\\ 2x=2\\ x=1\]

OpenStudy (yumeybitaa):

i got that one right , i did it on my own but i wanted to see if i was right thank you

OpenStudy (anonymous):

yw

OpenStudy (yumeybitaa):

that's all thank you for your help